OpenStudy (anonymous):

(x+y)^6 in the binomial theorem equation?

OpenStudy (anonymous):

\[\left(\begin{matrix}n \\ k-1\end{matrix}\right) x ^{n-(k-1)}y^{k-1}\] n=6 k=kth term in the sequence First one is